OpenStudy (anonymous):

If you were to solve the following system of equations by using a matrix, which of the following would be your coefficient matrix? 4x-3y=7 x+5y=13 (will post answer choices in a sec!) ***Please explain!! :) THanks! :D

OpenStudy (anonymous):

Answer choices A,B,C,D (top to bottom) :D

jimthompson5910 (jim_thompson5910):

4x-3y=7 x+5y=13 is the same as 4x-3y=7 1x+5y=13 your coefficients are the numbers that sit in front of the variables

OpenStudy (anonymous):

i'm confused though, does the matrix include the solution?? or just the other parts? (the equation)

jimthompson5910 (jim_thompson5910):

the coefficient matrix is only composed of the coefficients

jimthompson5910 (jim_thompson5910):

so what are the coefficients in this case?

OpenStudy (anonymous):

ohh so it's like this then?

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

so my answer is D? :)

jimthompson5910 (jim_thompson5910):

you got it, D is the answer

OpenStudy (anonymous):

yay! :) thanks!! :) And thanks for the medal!! haha :D

jimthompson5910 (jim_thompson5910):

np